3x - 1/2y = 5
1/3x + 1/4y = 3
Solve by elimination
Answer's are X=3 and Y=8

Answer by checkley77(12569)

3x - 1/2y = 5
1/3x + 1/4y = 3 MULTIPLY BY -9 & ADD.
-3X-9Y/4=-27
3X-Y/2=5
----------------------
-9Y/4-Y/2=-22
(-9Y-2Y)/4=-22
-11Y=-88
Y=-88/-11
Y=8 ANSWER.
3X-8/2=5
3X-4=5
3X=5+4
3X=9
X=9/3
X=3 ANSWER.
PROOF;
3/3+8/4=3
1+2=3
3=3